# FIPER: Generalizable Factorized Fields for Joint Image Compression and Super-Resolution

### v1

@@ 3.3 Super-Resolution with Factorized Fields
Our model comprises three main components: Coefficient Backbone, Basis Swin Transformer, and Factorized Fields Reconstruction. As shown in Fig. 2, the process begins with $I_{LR}\in\mathbb{R}^{3\times H\times W}$. The Coefficient Backbone extracts features $X_{\text{coeff}}\in\mathbb{R}^{C_{c}\times H_{c}\times W_{c}}$, which are then used to generate coefficients $c$ through convolution and pixel shuffle operations, and fed into the Basis Swin Transformer to produce a multi-scale basis $b=\{b_{1},...,b_{N}\},\ b_{i}\in\mathbb{R}^{C_{b_{i}}\times H_{b_{i}}\times W_{b_{i}}}$. The coefficients and basis are combined to reconstruct $I_{SR}\in\mathbb{R}^{3\times sH\times sW}$ using Eq. 8, where $s$ is the scale factor. In our experiments, $\gamma_{i}(x)$ is a Sawtooth function: $\gamma_{i}(x)=x\mod k_{i}$, with $k_{i}\in R$. We optimize model parameters using an $L_{1}$ loss function. To demonstrate our method’s effectiveness, we use existing SR methods chen2023activating, DAT, ATD as the Coefficient Backbone. For the Basis Swin Transformer, we employ Swin Transformer Blocks liu2022swinv2 with a series of downsampling operations. We use a dilation-like downsampling technique (Fig. 1(c)) to accommodate the sawtooth sampling pattern. The final basis is refined using additional upsampling and convolution layers.
@@ 3.4 Image Compression with Factorized Fields and Super-Resolution
Image Compression, at its core, is to strike a balance between the amount of information contained in the latent bits and the final image quality. However, as discussed in Sec. 1,  2, and 3.2, most recent works draw emphasis on how to retrieve implicit elements through designing different architectures, such as analysis transforms and entropy models koyuncu2024efficient, li2024groupedmixer, or decompression modules song2024wavelet-compression, duan2023cross-component, while this paper aims at addressing the representation itself to better capture the structural correlations and thus achieve better image quality through explicit modeling of frequency components and Factorized Fields formulation as in Eq. 7. In addition, with our trained SR model described in Sec. 3.3, it intuitively serves as a strong prior for information recovery, i.e., it contains extensive knowledge of how to reconstruct missing details and enhance image quality by leveraging learned patterns from the training data. Thus, since Super-Resolution and Image Compression share the core principle of reconstructing and enhancing image details from low-quality sources, we can effectively integrate this prior into the compression pipeline. In the following, we respectively present the proposed joint framework for the single and multi-image compression tasks.
@@ 3.4.1 Single Image Compression
The overall pipeline is shown in Fig. 3(b). To demonstrate the robustness of our representation and the effectiveness of the SR prior, the compression and decompression networks greatly follow imc_lic_tcm, with only the synthesis transform replaced by our SR pipeline, where the details can be referenced in Supplementary Materials. In practice, the training is performed in two stages. After we obtain the trained SR prior, the model is fine-tuned with a lower learning rate alongside the compression module, which is then trained end-to-end with the loss function defined in Eq. 5.
@@ 3.4.2 Multi-Image Compression
$b_{i}(h,w)=F_{\text{merge}}\left(\{b^{n}_{i}(h,w)\ |\ n\in{1,...,M}\}\right),\ 0\leq h<H_{b_{i}},\ 0\leq w<W_{b_{i}}.$ (9)
Here, $F_{\text{merge}}$ is a standard transformer, similar in architecture to that described in (dinov2). We treat the bases as tokens and prepend the sequence with a $CLS$ token, concluding the sequence to form the final merged basis.
@@
Table 1: Quantitative comparisons on 4×4$\times$4 × super-resolution with state-of-the-art methods. The best results are colored red. The models with †are those who use same-task pretraining chen2023activating. Please refer to quantitative results in Sec. 4.1 for details.
@@
SwinIR (swinir)(ICCV2021W) 28.01 119.68 3,826 32.93 0.9043 29.15 0.7958 27.95 0.7484 27.56 0.8273 32.22 0.9273
ATD (zhang2024transcending)(CVPR2024) 20.26 77.10 6,572 33.14 0.9061 29.25 0.7976 28.02 0.7524 28.22 0.8414 32.65 0.9308
DAT (DAT))(ICCV2023) 14.80 61.66 4,192 33.15 0.9062 29.29 0.7983 28.03 0.7518 27.99 0.8365 32.67 0.9301
RGT (RGT)(ICLR2024) 13.37 834.25 3,404 33.16 0.9066 29.28 0.7979 28.03 0.7520 28.09 0.8388 32.68 0.9303
HAT (chen2023hat)(CVPR2023) 20.77 86.02 3,692 33.18 0.9073 29.38 0.8001 28.05 0.7534 28.37 0.8447 32.87 0.9319
HAT-L (chen2023hat)(CVPR2023) 40.84 167.27 6,804 33.30 0.9083 29.47 0.8015 28.09 0.7551 28.60 0.8498 33.09 0.9335
@@ 4.1 Image Super-Resolution
We conduct extensive experiments to validate the effectiveness of our Factorized Fields representation for Super-Resolution tasks. Following the strategy outlined in (chen2023hat) and (drct), we adapt the same-task pretraining approach for all the Super-Resolution models. Unlike these previous works, we leverage the SA-1B dataset from (sam), which includes approximately 10 million images, where we randomly sample 4 million for training, which is much less that of ImageNet (imagenet), which is used for chen2023hat and drct pretraining. Note that we use SA-1B just for its content-rich and high-resolution images since SA-1B is designed for training segmentation models, whereas ImageNet focuses primarily on single-class prediction. We further conduct experiments to compare training performance with SA-1B and ImageNet in Tab. 1. While the performances are mostly equivalent, we witness faster convergence with SA-1B during training.
To be more focused on the representation itself, we utilize various pre-trained SR models (SwinIR (swinir), HAT (chen2023activating), DAT (DAT), and ATD (zhang2024transcending)) as the Coefficient Backbone, with a consistent Basis Transformer architecture. Training involves initializing the Coefficient Backbones from pre-trained SR models and randomly initializing the Basis Transformers. Models are pre-trained for 300k iterations on the SA-1B dataset. After pretraining, we use DF2K (DIV2K (div2k) + Flickr2K (flickr2k)) as the finetuning dataset following (swin2sr, chen2023hat) for 200k iterations.
For pretraining, we utilize AdamW optimizer with learning rate 1e-4, batch size 16, betas (0.9, 0.99), and other parameters set to PyTorch default, while we use learning rate 1e-5 during finetuning. Throughout training, the input is randomly cropped to $256\times 256$ and bicubicly resized to $64\times 64$ for Coefficient Backbone input. As for the hyperparameter of Factorized Fields, the number $N$ of coefficient and basis is set to 6, and the scalars $\alpha_{j}$ are set to {1, 4, 16, 64} since we want to capture both base frequency $\alpha_{j}=1$ and high frequency $\alpha_{j}=64$ information.
@@
Tab. 1 presents the quantitative comparison between our approach and state-of-the-art (SoTA) methods. We evaluate the methods using five benchmark datasets, including Set5 (set5), Set14 (set14), BSD100 (bsd100), Urban100 (urban100), and Manga109 (manga109). For quantitative metrics, PSNR and SSIM are reported. The average relative improvement of 204.4% in PSNR over the baseline across the five datasets is calculated using the formula $(c-b)/(a-b)$, where $b$ represents the PSNR of a representative baseline, SwinIR, $a$ represents the PSNR of SOTA, HAT-L, and $c$ represents the PSNR of our best model, HAT-L-F. This formula measures the relative performance gain of our model compared to the gap between HAT-L and SwinIR.
To validate the effectiveness of our framework, we employ three different SoTA SR models—ATD ATD, DAT DAT, and HAT chen2023activating—as the Coefficient Backbone in our pipeline, which we denote as ATD-F, DAT-F, and HAT-F, respectively. These models exhibit significant improvements when compared to their counterparts ATD-L, DAT-L, and HAT-L, which possess similar parameter counts. It is important to note that only chen2023activating provides a large-scale model. To maintain fairness, we scale up ATD and DAT to match this size and train them under the same configuration as our model, including both pretraining and fine-tuning stages. To ensure an equitable comparison with other methods, we further train another model, HAT-F-ImageNet, using ImageNet as the pretraining dataset, following the protocols outlined in IPT, chen2023activating. The results demonstrate that its performance remains consistent with only minor perturbations.
Furthermore, in traditional Fourier Series and other image processing methods jpeg, the basis is typically derived first and then used to compute the coefficients. In contrast, our method derives the coefficient features first, as illustrated in Fig. 2. To explore this difference, we develop another variant of our model, denoted HAT-F-Basis-First, where we reverse the order of operations. In this case, we first pass the image through the Basis Swin Transformer and then use the resulting basis features and the image input to derive the coefficients. This approach, however, leads to a gigantic performance drop, showing the importance of the order of the pipeline. Specifically, we argue that in our pipeline, the Coefficient Backbone functions more as a feature extraction module, where the refined features facilitate downstream basis extraction.
@@
{subfigure}
[b]0.32 Refer to caption {subfigure}[b]0.32 Refer to caption {subfigure}[b]0.32 Refer to caption
Figure 5: Kodak
Figure 6: CLIC
Figure 7: Tecnick
@@ 4.2 Single- and Multi-Image Compressions
We evaluate our Factorized Fields representation for image compression tasks, comparing it against state-of-the-art methods. Following our Super-Resolution setting in Experiment Setup from Sec. 4.1, we use the same set of SA-1B for training. To emphasize our representation, we initialize compression and decompression modules in Fig.3 from pre-trained imc_lic_tcm and the SR Module from those in Sec. 4.1, we then train the pipeline end to end on 256x256 patches for 200k iterations, with AdamW optimizer (loshchilov2017decoupled) with learning rate 1e-5, batch size 16, betas (0.9, 0.99) and other parameters set to PyTorch default.
We integrate our SR Module with the pre-trained compression module TCM, creating TCM-HAT-F and TCM-HAT-L-F models. TCM-HAT-F-multi represents the multi-image compression pipeline. For multi-image compression in Fig. 3(c), we set the Basis Swin Transformer and the Basis Merging Transformer to be trainable while the other parts remain frozen.
@@
We compare our model with State-of-the-Art learned end-to-end image compression algorithms, including (liu2023learned), (Chen2022), zou2022the, xie2021enhanced, cheng2020learned, ballé2018variational, li2024groupedmixer, jiang2023mlic, channel-autoregressive, bellard_bpg, qi2023hyperthumbnail, and imc_elic. The classical image compression codec, VVC (vtm), is also tested by using VTM12.1.The rate-distortion performance on various datasets, including Kodak, Tecnick’s old test set with resolution 1200$\times$1200, and CLIC Professional Validation, is shown in Fig. 8.
@@
Table 3: Comparison of improvements of Factorized Fields. $\psi$ and $\alpha$ are the same in Eq. 8
@@ 4.3 Ablation Studies
Influence of SR Priors in Image Compression.
We conduct experiments with various configurations to verify the proposed image compression pipeline’s effectiveness. As shown in Tab. 4, we present the quantitative performance of models trained with different values of $\lambda$ in Eq. 5. Specifically, TCM-HAT refers to substituting our SR Module with the original HAT chen2023activating in the pipeline illustrated in Fig. 3.b. TCM-HAT-F represents our complete pipeline, while TCM-HAT-F-Scratch denotes the same pipeline but with the SR Module initialized randomly. Our results demonstrate that integrating SR priors with image compression improves performance, and our proposed representation further enhances results. This highlights the robustness of our Factorized Fields in capturing fine details in image regression tasks.
Method Kodak CLIC Tecnick
[…]
TCMimc_lic_tcm 0.1533 30.0834 0.2983 32.5841 0.6253 36.1345 0.1214 31.8207 0.2235 34.2098 0.4503 37.1201 0.1268 32.0588 0.2193 34.3669 0.3981 36.9066

### v2

@@ 3.2 Super-Resolution with Factorized Fields
Our model comprises three main components: Coefficient Backbone, Basis Swin Transformer, and Factorized Fields Reconstruction. As shown in  Fig. 3, the process begins with $I_{LR}\in\mathbb{R}^{3\times H\times W}$. The Coefficient Backbone extracts features $X_{\text{coeff}}\in\mathbb{R}^{C_{c}\times H_{c}\times W_{c}}$, which are then used to generate coefficients $c$ through convolution and pixel shuffle operations, and fed into the Basis Swin Transformer to produce a multi-scale basis $b=\{b_{1},...,b_{N}\},\ b_{i}\in\mathbb{R}^{C_{b_{i}}\times H_{b_{i}}\times W_{b_{i}}}$. The coefficients and basis are combined to reconstruct $I_{SR}\in\mathbb{R}^{3\times sH\times sW}$ using  Eq. 4, where $s$ is the scale factor. In our experiments, $\gamma_{i}(x)$ is a Sawtooth function: $\gamma_{i}(x)=x\mod k_{i}$, with $k_{i}\in R$. We optimize model parameters using an $L_{1}$ loss function. To demonstrate our method’s effectiveness, we use existing SR methods [14, 15, 75] as the Coefficient Backbone. For the Basis Swin Transformer, we employ Swin Transformer Blocks [51] with a series of downsampling operations. We use a dilation-like downsampling technique( Fig. 2c) to accommodate the sawtooth sampling pattern. The final basis is refined using additional upsampling and convolution layers.
@@ 3.3 Image Compression with Factorized Fields and Super-Resolution
Image Compression, at its core, is to strike a balance between the amount of information contained in the latent bits and the final image quality. However, as discussed in  Sec. 1,  Sec. 2, and Sec. 3.1, most recent works draw emphasis on how to retrieve implicit elements through designing different architectures, such as analysis transforms and entropy models [38, 42], or decompression modules [63, 22], while this paper aims at addressing the representation itself to better capture the structural correlations and thus achieve better image quality through explicit modeling of frequency components and Factorized Fields formulation as in  Eq. 3. In addition, with our trained SR model described in  Sec. 3.2, it intuitively serves as a strong prior for information recovery, i.e., it contains extensive knowledge of how to reconstruct missing details and enhance image quality by leveraging learned patterns from the training data. Thus, since Super-Resolution and Image Compression share the core principle of reconstructing and enhancing image details from low-quality sources, we can effectively integrate this prior into the compression pipeline. In the following, we respectively present the proposed joint framework for the single and multi-image compression tasks.
@@ 3.3.1 Single Image Compression
The overall pipeline is shown in  Fig. 5b. To demonstrate the robustness of our representation and the effectiveness of the SR prior, the compression and decompression networks greatly follow [48], with only the synthesis transform replaced by our SR pipeline, where the details can be referenced in Supplementary Materials. In practice, the training is performed in two stages. After we obtain the trained SR prior, the model is fine-tuned with a lower learning rate alongside the compression module, which is then trained end-to-end with the loss function defined as
$L=R(\hat{y})+R(\hat{z})+\lambda\cdot D(x,\hat{x}),$ (5)
@@ 3.3.2 Multi-Image Compression
$b_{i}(h,w)=F_{\text{merge}}\left(\{b^{n}_{i}(h,w)\ |\ n\in{1,...,M}\}\right),$ (6)
[…]
@@
Table 1: Quantitative comparisons on 4×4$\times$4 × super-resolution with state-of-the-art methods. The best results are colored red. The models with †are those who use same-task pretraining [14]. Please refer to quantitative results in  Sec. 4.1 for details.
@@
SwinIR [44](ICCV2021W) 28.01 119.68 3,826 32.93 0.9043 29.15 0.7958 27.95 0.7484 27.56 0.8273 32.22 0.9273
ATD [76](CVPR2024) 20.26 77.10 6,572 33.14 0.9061 29.25 0.7976 28.02 0.7524 28.22 0.8414 32.65 0.9308
DAT [15])(ICCV2023) 14.80 61.66 4,192 33.15 0.9062 29.29 0.7983 28.03 0.7518 27.99 0.8365 32.67 0.9301
RGT [16](ICLR2024) 13.37 834.25 3,404 33.16 0.9066 29.28 0.7979 28.03 0.7520 28.09 0.8388 32.68 0.9303
HAT [13](CVPR2023) 20.77 86.02 3,692 33.18 0.9073 29.38 0.8001 28.05 0.7534 28.37 0.8447 32.87 0.9319
HAT-L [13](CVPR2023) 40.84 167.27 6,804 33.30 0.9083 29.47 0.8015 28.09 0.7551 28.60 0.8498 33.09 0.9335
@@ 4.1 Image Super-Resolution
We conduct extensive experiments to validate the effectiveness of our Factorized Fields representation for Super-Resolution tasks. Following the strategy outlined in [13] and [18], we adapt the same-task pretraining approach for all the Super-Resolution models. Unlike these previous works, we leverage the SA-1B dataset from [35], which includes approximately 10 million images, where we randomly sample 4 million for training, which is much less that of ImageNet [20], which is used for [13] and [18] pretraining. Note that we use SA-1B just for its content-rich and high-resolution images since SA-1B is designed for training segmentation models, whereas ImageNet focuses primarily on single-class prediction. We further conduct experiments to compare training performance with SA-1B and ImageNet in  Tab. 1. While the performances are mostly equivalent, we witness faster convergence with SA-1B during training.
To be more focused on the representation itself, we utilize various pre-trained SR models (SwinIR [44], HAT [14], DAT [15], and ATD [76]) as the Coefficient Backbone, with a consistent Basis Transformer architecture. Training involves initializing the Coefficient Backbones from pre-trained SR models and randomly initializing the Basis Transformers. Models are pre-trained for 300k iterations on the SA-1B dataset. After pretraining, we use DF2K (DIV2K [1] + Flickr2K [46]) as the finetuning dataset following [19, 13] for 200k iterations.
For pretraining, we utilize AdamW optimizer with learning rate 1e-4, batch size 16, betas (0.9, 0.99), and other parameters set to PyTorch default, while we use learning rate 1e-5 during finetuning. Throughout training, the input is randomly cropped to $256\times 256$ and bicubicly resized to $64\times 64$ for Coefficient Backbone input. As for the hyperparameter of Factorized Fields, the number $N$ of coefficient and basis is set to 6, and the scalars $\alpha_{j}$ are set to {1, 4, 16, 64} since we want to capture both base frequency $\alpha_{j}=1$ and high frequency $\alpha_{j}=64$ information. More implementation details can be found in supplementary materials.
@@
Tab. 1 presents the quantitative comparison between our approach and state-of-the-art (SoTA) methods. We evaluate the methods using five benchmark datasets, including Set5 [5], Set14 [74], BSD100 [55], Urban100 [31], and Manga109 [56]. For quantitative metrics, PSNR and SSIM are reported. The average relative improvement of 204.4% in PSNR over the baseline across the five datasets is calculated using the formula $(c-b)/(a-b)$, where $b$ represents the PSNR of a representative baseline, SwinIR, $a$ represents the PSNR of SOTA, HAT-L, and $c$ represents the PSNR of our best model, HAT-L-F. This formula measures the relative performance gain of our model compared to the gap between HAT-L and SwinIR.
To validate the effectiveness of our framework, we employ three different SoTA SR models—ATD [75], DAT [15], and HAT [14]—as the Coefficient Backbone in our pipeline, which we denote as ATD-F, DAT-F, and HAT-F, respectively. These models exhibit significant improvements when compared to their counterparts ATD-L, DAT-L, and HAT-L, which possess similar parameter counts. It is important to note that only [14] provides a large-scale model. To maintain fairness, we scale up ATD and DAT to match this size and train them under the same configuration as our model, including both pretraining and fine-tuning stages. To ensure an equitable comparison with other methods, we further train another model, HAT-F-ImageNet, using ImageNet as the pretraining dataset, following the protocols outlined in [12, 14]. The results demonstrate that its performance remains consistent with only minor perturbations.
Furthermore, in traditional Fourier Series and other image processing methods [69], the basis is typically derived first and then used to compute the coefficients. In contrast, our method derives the coefficient features first, as illustrated in  Fig. 3. To explore this difference, we develop another variant of our model, denoted HAT-F-Basis-First, where we reverse the order of operations. In this case, we first pass the image through the Basis Swin Transformer and then use the resulting basis features and the image input to derive the coefficients. This approach, however, leads to a gigantic performance drop, showing the importance of the order of the pipeline. Specifically, we argue that in our pipeline, the Coefficient Backbone functions more as a feature extraction module, where the refined features facilitate downstream basis extraction.
@@
Multi-frame SR.
Due to the space limit, we provide additional comparisons on multi-frame SR in the supplementary materials.
(a) Kodak
(b) CLIC
(c) Tecnick
@@
Method Kodak CLIC Tecnick
[…]
TCM[48] 0.1533 30.0834 0.2983 32.5841 0.6253 36.1345 0.1214 31.8207 0.2235 34.2098 0.4503 37.1201 0.1268 32.0588 0.2193 34.3669 0.3981 36.9066
@@ 4.2 Single- and Multi-Image Compressions
We evaluate our Factorized Fields representation for image compression tasks, comparing it against state-of-the-art methods. Following our Super-Resolution setting in Experiment Setup from  Sec. 4.1, we use the same set of SA-1B for training. To emphasize our representation, we initialize compression and decompression modules in Fig. 5 from pre-trained [48] and the SR Module from those in Sec. 4.1, we then train the pipeline end to end on 256x256 patches for 200k iterations, with AdamW optimizer [52] with learning rate 1e-5, batch size 16, betas (0.9, 0.99) and other parameters set to PyTorch default.
We integrate our SR Module with the pre-trained compression module TCM, creating TCM-HAT-F and TCM-HAT-L-F models. TCM-HAT-F-multi represents the multi-image compression pipeline. For multi-image compression in Fig. 5.c, we set the Basis Swin Transformer and the Basis Merging Transformer to be trainable while the other parts remain frozen.
@@
We compare our model with State-of-the-Art learned end-to-end image compression algorithms, including [49], [11], [82], [73], [17], [2], [42], [32], [57], [4], [61], and [30]. The classical image compression codec, VVC [67], is also tested by using VTM12.1.The rate-distortion performance on various datasets, including Kodak, Tecnick’s old test set with resolution 1200$\times$1200, and CLIC Professional Validation, is shown in Fig. 7.
@@
Table 4: Comparison of improvements of Factorized Fields. $\psi$ and $\alpha$ are the same in Eq. 4.
@@ 4.3 Ablation Studies
Influence of SR Priors in Image Compression.
We conduct experiments with various configurations to verify the proposed image compression pipeline’s effectiveness. As shown in Tab. 3, we present the quantitative performance of models trained with different values of $\lambda$ in Eq. 5. Specifically, TCM-HAT refers to substituting our SR Module with the original HAT [14] in the pipeline illustrated in Fig. 5b. TCM-HAT-F represents our complete pipeline, while TCM-HAT-F-Scratch denotes the same pipeline but with the SR Module initialized randomly. Our results demonstrate that integrating SR priors with image compression improves performance, and our proposed representation further enhances results. This highlights the robustness of our Factorized Fields in capturing fine details in image regression tasks.